# TexHOI: Reconstructing Textures of 3D Unknown Objects in Monocular Hand-Object Interaction Scenes

### v1

@@ 5.3 Evaluation Metrics
Since the objects manipulated in HO3D [hampali2020honnotate] dataset have their ground truth 3D models and textures in YCB [xiang2017posecnn], we use them for our quantitative evaluation. For each video sequence in HO3D, we subsample one frame out of every 10 frames, resulting in around $200$-$400$ frames per sequence as our experimental data. We use Mitsuba [Mitsuba3] renderer and render the ground truth 3D models and textures to generate ground truth images corresponding to the poses of those sampled image frames, to compare our TexHOI’s predicted albedo with other baseline methods. The quantitative evaluation metrics used are as follows:
@@ 5.4 Baseline Comparisons
From Fig. 7, it is evident that while HOLD produces final images that closely resemble the input image, it achieves this by baking in the hand shadows and environmental reflections directly into the object’s texture. This, in turn, reduces the model’s ability to generalize across different lighting environments and hand poses. In contrast, our method, TexHOI, can disentangle the object’s intrinsic texture (albedo) from external environmental influences, resulting in a cleaner and more accurate reconstruction of the object’s surface. This observation is further supported by the evaluation metrics in Tab. 1. HOLD achieves better performance on metrics such as RGB similarity because it embeds hand shadows and environmental reflections into the predicted texture, making them appear closer to the ground truth in terms of pixel-wise accuracy. However, our method excels in producing albedo textures that are more accurate representations of the object’s inherent properties, without conflating them with transient environmental elements.
@@ 5.5 Ablation Studies
To evaluate the importance of different components in our approach, we conduct ablation studies by removing specific features, such as indirect illumination and hand occlusion handling. Specifically, we compare the results of our full methodology against versions without indirect illumination modeling and without explicit hand occlusion consideration. Our results, from Fig. 8 and Tab. 2 indicate that indirect illumination significantly contributes to the realism of the reconstructed texture, particularly in complex scenes where light bounces off the occluding hand surface. Without this component, the reconstructed textures appear flatter and lack depth. Similarly, ignoring hand occlusions results in visual artifacts, as the shadowing and occlusion effects caused by interacting hands are not properly accounted for, leading to inconsistencies in the texture and geometry. These experiments highlight the necessity of modeling both indirect illumination and occlusions to achieve high-quality, realistic reconstructions. The few examples where the version without indirect illumination surpasses our full model in Tab. 2 are again because the final prediction bakes the hand-environment interactions in the final prediction.
@@ 6 Conclusion and Future Work
Despite these advancements, there are several areas for future improvement. One potential direction is incorporating a more sophisticated model for indirect illumination, particularly one that accounts for the specific properties of skin and other materials under different lighting conditions, could enhance the realism of the reconstructions. Another promising direction is exploring temporal consistency across frames in dynamic sequences, avoiding minor temporal artifacts. By incorporating temporal coherence into the reconstruction process, we could achieve smoother transitions and more stable textures in video sequences. In future, the work can also be optimized end-to-end by incorporating composite surface-based rendering. With this end-to-end optimization, we do not have to perform 2-step optimization to first obtain corrected object and hand poses, and then optimize the surface texture. All these elements can be simultaneously optimized. Finally, extending our model to handle a broad range of interaction scenarios, such as multiple hands, or handling inter-penetrations of hand and object using collision detection or physical simulation modules, would make it more applicable to real-world use cases with complex interactions.

### v2

@@ 5.3 Evaluation Metrics
Since the objects manipulated in HO3D [21] dataset have their ground truth 3D models and textures in YCB [57], we use them for our quantitative evaluation. For each video sequence in HO3D, we subsample one frame out of every 10 frames, resulting in around $200$-$400$ frames per sequence as our experimental data. We use Mitsuba [1] renderer and render the ground truth 3D models and textures to generate ground truth images corresponding to the poses of those sampled image frames, to compare our TexHOI’s predicted albedo with other baseline methods. The quantitative evaluation metrics used are as follows:
@@ 5.4 Baseline Comparisons
From Fig. 7, it is evident that while HOLD produces final images that closely resemble the input image, it achieves this by baking in the hand shadows and environmental reflections directly into the object’s texture. This, in turn, reduces the model’s ability to generalize across different lighting environments and hand poses. In contrast, our method, TexHOI, can disentangle the object’s intrinsic texture (albedo) from external environmental influences, resulting in a cleaner and more accurate reconstruction of the object’s surface. This observation is further supported by the evaluation metrics in Tab. I. HOLD achieves better performance on metrics such as RGB similarity because it embeds hand shadows and environmental reflections into the predicted texture, making them appear closer to the ground truth in terms of pixel-wise accuracy. However, our method excels in producing albedo textures that are more accurate representations of the object’s inherent properties, without conflating them with transient environmental elements.
@@ 5.5 Ablation Studies
To evaluate the importance of different components in our approach, we conduct ablation studies by removing specific features, such as indirect illumination and hand occlusion handling. Specifically, we compare the results of our full methodology against versions without indirect illumination modeling and without explicit hand occlusion consideration. Our results, from Fig. 8 and Tab. II indicate that indirect illumination significantly contributes to the realism of the reconstructed texture, particularly in complex scenes where light bounces off the occluding hand surface. Without this component, the reconstructed textures appear flatter and lack depth. Similarly, ignoring hand occlusions results in visual artifacts, as the shadowing and occlusion effects caused by interacting hands are not properly accounted for, leading to inconsistencies in the texture and geometry. These experiments highlight the necessity of modeling both indirect illumination and occlusions to achieve high-quality, realistic reconstructions. The few examples where the version without indirect illumination surpasses our full model in Tab. II are again because the final prediction bakes the hand-environment interactions in the final prediction.
@@ 6 Conclusion and Future Work
Despite these advancements, there are several areas for future improvement. One potential direction is incorporating a more sophisticated model for indirect illumination, particularly one that accounts for the specific properties of skin and other materials under different lighting conditions, could enhance the realism of the reconstructions. Although the centers of the hand spheres remain outside the object’s surface, there can be inter-penetrations of the spheres with object near the fingertips. This causes negligible errors in the computation of hand occlusion, and can be an area of improvement in the future. Another promising direction is exploring temporal consistency across frames in dynamic sequences, avoiding minor temporal artifacts. By incorporating temporal coherence into the reconstruction process, we could achieve smoother transitions and more stable textures in video sequences. 3D Gaussian splat-based inverse rendering have not been explored in the area of hand-object interactions, according to our knowledge, and can be a promising direction of progress in the research. In the future, the work can also be optimized end-to-end using composite surface-based inverse rendering or a more efficient 3D Gaussian splat-based inverse rendering, eliminating the need for a 2-step process to correct poses and optimize texture, as all elements can be optimized simultaneously. Finally, extending our model to handle a broad range of interaction scenarios, such as multiple hands, or handling inter-penetrations of hand and object using collision detection or physical simulation modules, would make it more applicable to real-world use cases with complex interactions.